# Geometry and factorization of multivariate Markov chains with applications to the swapping algorithm

### v4

@@ 3.2 Generalization to $d$-temperature swapping algorithm with $d\geq 2$
The discussion so far in this section can be generalized to design a projection sampler based on the $d$-temperature swapping algorithm with $\mathbb{N}\ni d\geq 2$. The main results of this section (Corollary 3.4 and 3.5 below) state that various mixing time parameters of the leave-$\{1\}$-out chain based on a $d$-temperature swapping algorithm can be improved by at least a factor of the dimension of the underlying state space times the number of temperatures $d$. We also offer an intuitive explanation on the acceleration effect in Remark 3.1.
@@
In the terminologies of Jerrum et al. (2004), the projection chain (i.e. the notation $\overline{P}$ therein) of $P_{sw}$ with respect to the above partition is $P_{sw}^{(-1)}$, while the restriction chains (i.e. the notation $P_{i}$ therein) on the state space $\Omega_{x^{(-1)}}$ can be shown to be, for each $x^{(-1)}\in\mathcal{X}^{d-1}$,
@@
$\gamma((P_{sw})_{x^{(-1)}})=\dfrac{1}{2d}\gamma(P_{0}),\quad\alpha((P_{sw})_{x^{(-1)}})=\dfrac{1}{2d}\alpha(P_{0}).$ (16)
@@
$\gamma((P_{sw})_{x^{(-1)}})=\dfrac{1}{d}\dfrac{1}{N+1},\quad\alpha((P_{sw})_{x^{(-1)}})=\dfrac{1}{2d}\dfrac{1}{N+1}.$ (17)
We now recall the parameter $\Gamma$ (i.e. the notation $\gamma$ in Jerrum et al. (2004)) introduced earlier in (14). Let $x^{*}=\operatorname*{arg\,max}\mathcal{H}(x)$ and $y^{*}$ be chosen such that $y^{*}\neq x^{*}$. Let $\mathbf{x}^{*}\in\mathcal{X}^{d-1}$ be a $(d-1)$-dimensional vector with all entries equal to $x^{*}$. It can then readily be seen that
@@
Using again Theorem $1$ and Theorem $4$ in Jerrum et al. (2004) leads to
@@
In simulating $P_{sw}^{(-1)}$, we first sample according to the stationary distribution of the first coordinate $\pi_{0}$, followed by a step in the swapping algorithm. As the first coordinate is at stationarity, the swapping algorithm only needs to equilibrate the remaining $d-1$ coordinates.
@@
In the literature, it is noted in Madras and Zheng (2003) that a common choice is to set $d$ proportional to $N$. As a result, if we choose $d=N$, then the leave-$\{1\}$-out projection chain enjoys at least a speedup of a multiplicative factor of $N^{2}/\log(N)$ when compared with the original swapping algorithm in terms of the worst-case $L^{2}$ mixing time.
@@
Acknowledgements
Michael Choi is grateful for helpful discussions on large deviations with Pierre Del Moral, and on information theory with Cheuk Ting Li and Lei Yu. Michael Choi acknowledges the financial support of the project “MAPLE: Mechanistic Accelerated Prediction of Protein Secondary Structure via LangEvin Monte Carlo” with grant number 22-5715-P0001 under the NUS Faculty of Science Ministry of Education Tier 1 grant Data for Science and Science for Data collaborative scheme, project NUSREC-HPC-00001 and NUSREC-CLD-00001 for NUS HPC-AI Priority Projects for Research Program, as well as the startup funding of the National University of Singapore with grant number A-0000178-01-00. Youjia Wang gratefully acknowledges the financial support from National University of Singapore via the Presidential Graduate Fellowship. GW expresses gratitude to Shun Watanabe for insightful discussions and for highlighting references, specifically (Amari and Han, 1989) and (Watanabe, 2017). The work of GW was supported in part by the Special Postdoctoral Researcher Program (SPDR) of RIKEN and by the Japan Society for the Promotion of Science KAKENHI under Grant 23K13024.

### v5

@@ 3.2 Generalization to $d$-temperature swapping algorithm with $d\geq 2$
The discussion so far in this section can be generalized to design a projection sampler based on the $d$-temperature swapping algorithm with $\mathbb{N}\ni d\geq 2$. The main results of this section (Corollary 3.4 and 3.6 below) state that various mixing time parameters of the leave-$\{1\}$-out chain based on a $d$-temperature swapping algorithm can be improved by at least a factor of the dimension of the underlying state space times the number of temperatures $d$. We also offer an intuitive explanation on the acceleration effect in Remark 3.5.
@@
In the terminologies of [23], the projection chain (i.e. the notation $\overline{P}$ therein) of $P_{sw}$ with respect to the above partition is $P_{sw}^{(-1)}$, while the restriction chains (i.e. the notation $P_{i}$ therein) on the state space $\Omega_{x^{(-1)}}$ can be shown to be, for each $x^{(-1)}\in\mathcal{X}^{d-1}$,
@@
(17) $\gamma((P_{sw})_{x^{(-1)}})=\dfrac{1}{2d}\gamma(P_{0}),\quad\alpha((P_{sw})_{x^{(-1)}})=\dfrac{1}{2d}\alpha(P_{0}).$
@@
(18) $\gamma((P_{sw})_{x^{(-1)}})=\dfrac{1}{d}\dfrac{1}{N+1},\quad\alpha((P_{sw})_{x^{(-1)}})=\dfrac{1}{2d}\dfrac{1}{N+1}.$
We now recall the parameter $\Gamma$ (i.e. the notation $\gamma$ in [23]) introduced earlier in (15). Let $x^{*}=\operatorname*{arg\,max}\mathcal{H}(x)$ and $y^{*}$ be chosen such that $y^{*}\neq x^{*}$. Let $\mathbf{x}^{*}\in\mathcal{X}^{d-1}$ be a (d1)(d-1)-dimensional vector with all entries equal to $x^{*}$. It can then readily be seen that
@@
Using again Theorem $1$ and Theorem $4$ in [23] leads to
@@
In simulating $P_{sw}^{(-1)}$, we first sample according to the stationary distribution of the first coordinate $\pi_{0}$, followed by a step in the swapping algorithm. As the first coordinate is at stationarity, the swapping algorithm only needs to equilibrate the remaining d1d-1 coordinates.
@@
In the literature, it is noted in [29] that a common choice is to set $d$ proportional to $N$. As a result, if we choose $d=N$, then the leave-$\{1\}$-out projection chain enjoys at least a speedup of a multiplicative factor of $N^{2}/\log(N)$ when compared with the original swapping algorithm in terms of the worst-case $L^{2}$ mixing time.
@@
3.3 Numerical experiments
In this section, we present a simple bimodal example as in Section 1.1.1, where the last coordinate of $P_{sw}^{(-1)}$ mixes well while that of $P_{sw}$ is stuck at the region around one mode.
In view of the setting and notations in this section and Section 1.1.1, suppose the state space is $\mathcal{X}=\llbracket-n,n\rrbracket$ on which the target distribution is, for $x\in\mathcal{X}$,
$\pi_{\beta}(x)\propto 2^{|x|}=e^{-\beta\mathcal{H}(x)},$
where we take $\beta=\ln 2$ and $\mathcal{H}(x)=-|x|$. There are two modes of this distribution at $\pm n$ respectively. In this context, $\pi_{0}$ is simply the discrete uniform distribution on $\mathcal{X}$ which can be sampled at ease.
For reproducibility, the code used in our experiments is available at https://github.com/mchchoi/factorization/tree/main. We now state the parameters used in the experiments:
  • $n=100$.
  • $d=3$ with temperature ladder $(\beta_{1},\beta_{2},\beta_{3})=(0,\frac{\ln 2}{2},\ln 2)$.
  • The proposal chain of Metropolis-Hastings during the level move of the swapping algorithm moves from $x$ to $\min\{x+1,n\}$ and $\max\{x-1,-n\}$ with probability $1/2$, and 0 otherwise.
  • All samplers $P_{sw}^{(-1)},P_{sw}$ are initialized at 100-100, the mode on the left, and are simulated for $100,000$ steps.
The results are summarized and presented in Figure 4, Table 3 and 4.
First, we note that $P_{sw}$ does not exhibit mixing: from the traceplot, histogram and empirical mean, it only explores the basin around the left mode at 100-100 and do not traverse to the right mode at $100$ in the experiment.
Second, from the traceplots and histograms we see that $P_{sw}^{(-1)}$ is able to hop between the two modes effectively. From Table 1 the empirical distribution generated by $P_{sw}^{(-1)}$ is notably closer to the ground truth $\pi_{\beta}$ than that generated by $P_{sw}$. From Table 4 the empirical mean and second moment generated by $P_{sw}^{(-1)}$ are also closer to the respective ground truth than that generated by $P_{sw}$. These results give empirical evidence that it is advantageous to use the projection sampler $P_{sw}^{(1)}$ over $P_{sw}$ to sample from $\pi_{\beta}$.
(a) Traceplot and histogram of the trajectories of the last coordinate of $P_{sw}$.
(b) Traceplot and histogram of the trajectories of the last coordinate of $P_{sw}^{(-1)}$.
Figure 4: Numerical experiments comparing the two samplers $P_{sw},P_{sw}^{(-1)}$ with target distribution being the V-shaped $\pi_{\beta}(x)\propto 2^{|x|}$.
Sampler $\widetilde{D}_{TV}(\widehat{\pi}_{\beta},\pi_{\beta})$ $\widetilde{D}_{KL}(\widehat{\pi}_{\beta}\|\pi_{\beta})$
$P_{sw}$ (last coordinate) 0.50 0.69
$P_{sw}^{(-1)}$ (last coordinate) 0.01 0.00
Table 3: Comparison of total variation distance and KL divergence between $\widehat{\pi}_{\beta}$ and the ground truth $\pi_{\beta}$, where $\widehat{\pi}_{\beta}$ is the empirical distribution formed by the trajectories of the samplers.
Sampler Mean Second moment
$P_{sw}$ (last coordinate) -99.01 9804.85
$P_{sw}^{(-1)}$ (last coordinate) -2.01 9804.27
Truth $\pi^{(1)}$ 0 9803.00
Table 4: Comparison of the first and second moment between the samplers and the ground truth $\pi_{\beta}$.
Acknowledgments
The authors gratefully acknowledge the constructive comments from the associate editor and two reviewers. Michael Choi is grateful for helpful discussions on large deviations with Pierre Del Moral, and on information theory with Cheuk Ting Li and Lei Yu. Michael Choi acknowledges the financial support of the projects A-8001061-00-00, NUSREC-HPC-00001, NUSREC-CLD-00001, A-0000178-01-00, A-0000178-02-00 and A-8003574-00-00 at National University of Singapore. Youjia Wang gratefully acknowledges the financial support from National University of Singapore via the Presidential Graduate Fellowship. GW expresses gratitude to Shun Watanabe for insightful discussions. The work of GW was supported in part by the Special Postdoctoral Researcher Program (SPDR) of RIKEN and by the Japan Society for the Promotion of Science KAKENHI under Grant 23K13024.